OpenStudy (anonymous):

Given f(x)=-x2+2x, find f(-3)

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

Some one please help!

OpenStudy (shamil98):

x^2 + 2x when f(3) 3^2 + 2(3) = 9 + 6 = 15

OpenStudy (anonymous):

so its 15 then?

OpenStudy (shamil98):

yes

OpenStudy (anonymous):

alright thank you!

OpenStudy (anonymous):

can you help with a few more?